The length of a rectangle is 7 times its width. If its perimeter is 120 cm, what is the area?
Answer by scott8148(6628)

let x=width, so 7x=length __ 120=2(x)+2(7x) __ 120=16x __ 7.5=x

area equals length times width __ a=52.5*7.5